# UniTracker: Learning Universal Whole-Body Motion Tracker for Humanoid Robots

### v1

@@ Abstract
Humanoid robots must achieve diverse, robust, and generalizable whole-body control to operate effectively in complex, human-centric environments. However, existing methods, particularly those based on teacher-student frameworks often suffer from a loss of motion diversity during policy distillation and exhibit limited generalization to unseen behaviors. In this work, we present UniTracker, a simplified yet powerful framework that integrates a Conditional Variational Autoencoder (CVAE) into the student policy to explicitly model the latent diversity of human motion. By leveraging a learned CVAE prior, our method enables the student to retain expressive motion characteristics while improving robustness and adaptability under partial observations. The result is a single policy capable of tracking a wide spectrum of whole-body motions with high fidelity and stability. Comprehensive experiments in both simulation and real-world deployments demonstrate that UniTracker significantly outperforms MLP-based DAgger baselines in motion quality, generalization to unseen references, and deployment robustness, offering a practical and scalable solution for expressive humanoid control.
@@
Figure 1: We deploy our UniTracker on a real humanoid robot, enabling it to perform a diverse range of motions, including (1)stretching, (2)traditional Chinese martial arts, (3)dancing, (4)high kicks, (5)ball-kicking, (6)squatting, and (7)various other common motions.
I INTRODUCTION
@@
Humanoid robots have garnered growing interest in the robotics community for their human-like morphology, which equips them with the potential to perform a wide range of tasks traditionally carried out by humans. To function effectively in real-world, human-centric environments, these robots must exhibit not only physical versatility but also robust and expressive motor control. Among the key enablers of such capabilities is whole-body control, which coordinates multiple joints and limbs to perform complex tasks while ensuring stability, adaptability, and energy efficiency.
Within this broader category, whole-body motion tracking plays a crucial role. It involves reproducing rich human motions, such as walking, kicking, turning, and dancing, based on reference trajectories from demonstrations or motion datasets. Unlike general task-driven control, tracking requires precise alignment with reference poses while maintaining robustness under real-world constraints such as sensor noise, partial observability, and actuation delay. This makes the problem especially challenging, as it demands both fidelity and stability in high-dimensional action spaces.
@@
Recent research has explored various control interfaces tailored to whole-body motion tracking. These can be broadly categorized into dense and sparse control signals. Dense signals, such as teleoperation [ze2025twistteleoperatedwholebodyimitation, lu2025mobiletelevisionpredictivemotionpriors, fu2024humanplushumanoidshadowingimitation, he2024learninghumantohumanoidrealtimewholebody, qiu2025humanoidpolicyhuman, he2024omnih2ouniversaldexteroushumantohumanoid], offline motion datasets [shao2025langwbclanguagedirectedhumanoidwholebody, shi2025adversariallocomotionmotionimitation, he2025asapaligningsimulationrealworld, ji2025exbody2advancedexpressivehumanoid, cheng2024expressivewholebodycontrolhumanoid, chen2025gmt], and video-based motion estimation [mao2024learningmassivehumanvideos, allshire2025visualimitationenablescontextual], provide rich trajectory-level information. In contrast, sparse signals such as high-level task commands and VR-based guidance [he2025hoverversatileneuralwholebody, xue2025unifiedgeneralhumanoidwholebody] offer minimal information and often lead to reduced motion quality. In this work, we focus on universal whole-body tracking, where the input is a reference motion sequence and the goal is to track it robustly and expressively using a single policy.
A widely adopted approach is the teacher-student framework, where a teacher policy is trained with privileged observations to track the reference motion with high accuracy. The student policy then learns to mimic the teacher using only deployable observations. While effective in principle, existing student training methods—particularly those based on standard DAgger [ross2011reductionimitationlearningstructured] often converge to averaged behaviors, losing the richness of the original motions. Moreover, these methods struggle to generalize to unseen motion sequences.
To address this limitation, we propose UniTracker, a unified and expressive whole-body tracking framework that introduces a Conditional Variational Autoencoder (CVAE) into the student policy. By explicitly modeling a structured latent space conditioned on future motion references, the policy learns to generate diverse and high-fidelity behaviors even under partial observations. From a probabilistic perspective, the latent variable captures ambiguity in the mapping from observations to actions, enabling the policy to approximate a distribution over plausible motor behaviors rather than collapsing to a single deterministic output.
Unlike prior works that directly include reference signals into the policy input, our approach encourages semantic abstraction by guiding behavior generation through latent intention modeling. We further demonstrate that design choices such as residual KL formulation, latent dimension, and future window size significantly influence both training stability and downstream performance. Our training dataset is derived from the PHC-filtered AMASS corpus, which removes overly aggressive motions and improves real-world policy viability.
UniTracker is evaluated extensively in both simulation and real-world settings. Experiments on a 29-DoF Unitree G1 humanoid show that our policy can track 8k motions, including highly dynamic ones with a single network. Compared to strong teacher-student baselines without CVAE, UniTracker achieves better tracking accuracy, motion smoothness, and generalization. Furthermore, we demonstrate its application to text-driven motion generation and video-based motion estimation, where the policy successfully executes motions generated from external inputs not seen during training.
Our main contributions are summarized as follows: i)Diversity-aware policy via CVAE modeling: We introduce a Conditional Variational Autoencoder (CVAE) into the student policy to explicitly capture the distributional characteristics of human motions, allowing the policy to generate diverse and expressive behaviors under partial observations. ii)Improved generalization to unseen motions: By modeling motion diversity in a structured latent space, our approach enables the policy to generalize effectively to out-of-distribution motions, outperforming conventional DAgger-based methods. iii)Robust real-world tracking with a single universal policy: We demonstrate that our approach enables a humanoid robot to robustly track thousands of diverse human motions using a single policy, with successful deployment in both simulation and the real world.
@@
Figure 2: An overview of UniTracker: In Stage 1, we train a teacher policy using oracle states via goal-conditioned reinforcement learning. In Stage 2, we distill the policy into a deployable form using a CVAE-based DAgger framework. In Stage 3, we deploy the policy using the CVAE prior and decoder for execution under partial observations. The training dataset is derived from the AMASS dataset, filtered by PHC to remove physically infeasible motions.
II METHOD
II-A Problem Formulation
@@
We formulate the problem of humanoid robot whole-body motion tracking as a goal-conditioned reinforcement learning (RL) task, where a policy $\pi$ is trained to track reference motions at the whole-body level. The state $s_{t}$ comprises the robot’s proprioceptive information $s_{t}^{p}$ and the goal $s_{t}^{g}$ which specifies the target state for all body parts. The reward function $r_{t}=R(s_{t}^{p},s_{t}^{g})$, defined in terms of the agent’s proprioception and goal state, yields dense signals to guide policy optimization. To better focus on motion tracking at the whole-body level, we fix the wrist joints of our 29-degree-of-freedom (DoF) Unitree G1 robot[unitreeg1], reducing the action space to 23 dimensions. The action $a_{t}\in\mathbb{R}^{23}$ specifies target joint positions, which are executed via a PD controller to actuate the robot. For policy optimization, we employ Proximal Policy Optimization (PPO) [schulman2017proximalpolicyoptimizationalgorithms] to maximize the expected cumulative discounted reward $\mathbb{E}\left[\sum_{t=1}^{T}\gamma^{t-1}r_{t}\right]$.
This setup raises two critical questions in developing a universal yet robust humanoid robot motion tracking policy: 1) how to maximize the policy’s expressiveness to track a diverse range of motions and 2) how to minimize the sim-to-real performance degradation caused by the dynamic mismatch between simulation and the real world.
UniTracker addresses these challenges through a CVAE-based two stage training pipeline. First, we train a perturbation-free oracle policy in simulation using filtered reference motions retargeted from large-scale human motion dataset. Second, we distill the oracle policy into a hierarchical controller via a Conditional Variational Autoencoder (CVAE).The resulting universal whole-body motion tracker (UniTracker) is shown in Figure 2.
The remainder of this section is structured as follows. First, we describe the curation of a high-quality, feasible humanoid robot motion dataset for policy training in Section II-B and then detail the training of the oracle policy in simulation which attempts to maximize the expressiveness of our motion tracker in Section II-C. Further we explain the distillation of the oracle policy using CVAE in Section II-D
@@
Train from Scratch 58.32 145.59 12.35 10.11 139.77 11.92 10.04
@@
TABLE I: Baselines Comparison and Ablation Studies We compare our method with two baseline approaches and observe that it consistently outperforms both. Additionally, we conduct ablation studies on architectural design, KL residuals, KL coefficient, future window size, and latent dimensionality. Finally, we analyze the diversity of CVAE comparing the stochastic latent and the deterministic latent. The final configuration is selected based on the best overall performance observed across these settings.
II-B Humanoid Motion Dataset Curation
A large-scale humanoid motion dataset serves as fuel for training a universal motion tracker. Our dataset is primarily derived from the publicly available AMASS [mahmood2019amassarchivemotioncapture] dataset, filtered to exclude interactions and sequences shorter than 10 frames. This yields a final training set of 11,313 human motions, represented using the SMPL [SMPL:2015] parameters. The SMPL model parameterizes the human body through shape parameters $\beta\in\mathbb{R}^{10}$, pose parameters $\theta\in\mathbb{R}^{24\times 3}$ and root translation $p\in\mathbb{R}^{3}$. $S$ denotes the SMPL function, where $S(\beta,\theta,p):\beta,\theta,p\rightarrow\mathbb{R}^{6980\times 3}$ maps the SMPL parameters to the positions of vertices of a triangular human mesh.
To bridge the embodiment gap between SMPL human model and humanoid robots, we employ a two-stage retargeting approach inspired by H2O [he2024learninghumantohumanoidrealtimewholebody]. First, we carefully select 16 corresponding body links and optimize the shape parameter $\beta^{\prime}$ for humanoid robots by minimizing the distances between selected links in the rest pose. Second, leveraging the optimized $\beta^{\prime}$ alongside the original pose $\theta$ and translation $p$ from dataset, we perform gradient descent over the humanoid robot’s root translation, root orientation and joint positions to minimize the distances between selected links throughout the whole sequence. Additional regularization terms are added to avoid aggressive behaviors and ensure temporal smoothness.
II-C Oracle Policy Training in Simulation
Reward Design. We formulate the reward $r_{t}$ as a weighted sum of three components: 1) task rewards for motion tracking, 2)regularization, and 3) penalty, detailed in Table III. We apply curriculum learning to the regularization terms and penalty terms such that the policy could better focus on the motion tracking task itself and gradually take the penalty and regularization into account for more reasonable behaviors.
Early Termination and Reference State Initialization. In the early stage of training, the agent is prone to falling, resulting in the collection of invalid data that hinders effective learning. To address this issue, we follow prior works [Peng_2018, luo2023perpetualhumanoidcontrolrealtime] and introduce two early termination conditions: 1) orientation: the projected gravity on x or y axis exceeds 0.8; 2) tracking tolerance: the average link distance between the robot and reference motions is further than 0.5m. Proper task initialization is also crucial for RL training. We employ the Reference State Initialization [Peng_2018, he2025asapaligningsimulationrealworld] framework, where the starting point of the reference motion is randomly sampled for the policy to track. The robot’s initial state, including the root position, orientation, linear and angular velocities, as well as joint positions and velocities, is then derived from the corresponding reference pose. This initialization strategy substantially enhances motion tracking training by enabling the policy to learn different motions phases in parallel, rather being constrained to a sequential learning process.
Domain Randomization. Domain randomization has been proved to be a critical source of robustness and generalization for successful sim-to-real transfer [cheng2024expressivewholebodycontrolhumanoid, he2024learninghumantohumanoidrealtimewholebody, he2024omnih2ouniversaldexteroushumantohumanoid]. Existing domain randomization terms can be broadly categorized into two classes: asset property randomization and environmental dynamic randomization. Asset property randomization, focusing on parameters such as friction coefficients, center of mass offsets, and link masses, aims to prevent the policy from overfitting to a specific asset configuration. In contrast, environmental dynamic randomization, focusing on PD gains, applied torques, or other external perturbations (e.g., pushing the robot), introduces dynamic variations to simulate the environmental uncertainties encountered during real-world deployment. While previous works typically apply both forms of randomization concurrently throughout training, we propose decoupling them to better balance the tradeoff between motion tracking expressiveness and robustness. Hence, in the first stage, we exclusively randomize asset properties to maximize the policy’s ability to accurately track diverse motions. This approach not only builds an upper bound for subsequent policies but also yields a proxy agent capable of generating high-quality action signals in simulators.
Figure 3: The generalization capability of UniTracker in MuJoCo
II-D Hierarchical Controller via Online Distillation
$\log$ $P(a_{t}|s_{t}^{p-d},s_{t}^{g-d})\geq E_{\varepsilon(z|s_{t}^{p-o},s_{t}^{g-o})}[\log D(a_{t}|s_{t}^{p-d},z_{t})$
$\rho(z_{t}|s_{t}^{p-d},s_{t}^{g-d})=\textit{N}$ $(\mu^{\rho}(s_{t}^{p-d},s_{t}^{g-d}),\sigma^{\rho}(s_{t}^{p-d},s_{t}^{g-d}))$
$\sigma^{\varepsilon}(s_{t}^{p-o},s_{t}^{g-o}))$
$L=L_{action}+\beta L_{KL}$
where $L_{action}=||a_{t}^{deploy}-a_{t}^{oracle}||_{2}^{2}$ and $L_{KL}$ is the KL divergence between the prior and encoder.
III EXPERIMENT
III-A Experiment Setup

### v2

@@ Abstract
Achieving expressive and generalizable whole-body motion control is essential for deploying humanoid robots in real-world environments. In this work, we propose UniTracker, a three-stage training framework that enables robust and scalable motion tracking across a wide range of human behaviors. In the first stage, we train a teacher policy with privileged observations to generate high-quality actions. In the second stage, we introduce a Conditional Variational Autoencoder (CVAE) to model a universal student policy that can be deployed directly on real hardware. The CVAE structure allows the policy to learn a global latent representation of motion, enhancing generalization to unseen behaviors and addressing the limitations of standard MLP-based policies under partial observations. Unlike pure MLPs that suffer from drift in global attributes like orientation, our CVAE-student policy incorporates global intent during training by aligning a partial-observation prior to the full-observation encoder. In the third stage, we introduce a fast adaptation module that fine-tunes the universal policy on harder motion sequences that are difficult to track directly. This adaptation can be performed both for single sequences and in batch mode, further showcasing the flexibility and scalability of our approach. We evaluate UniTracker in both simulation and real-world settings using a Unitree G1 humanoid, demonstrating strong performance in motion diversity, tracking accuracy, and deployment robustness.
@@
Figure 1: We deploy our UniTracker on a real humanoid robot, enabling it to perform a diverse range of motions, including (1)stretching, (2)traditional Chinese martial arts, (3)dancing, (4)high kicks, (5)ball-kicking, (6)various other common motions, and (7)challenging motions by fast adaption.
Keywords: Whole-Body Humanoid Robot Tracker, Imitation Learning, Reinforcement Learning
1 Introduction
@@
Humanoid robots have garnered growing interest in the robotics community for their human-like morphology, which equips them with the potential to perform a wide range of tasks traditionally carried out by humans. To function effectively in real-world, human-centric environments, these robots must exhibit not only physical versatility but also robust and expressive motor control. Among the key enablers of such capabilities is whole-body control, which coordinates multiple joints and limbs to perform complex tasks while ensuring stability, expressiveness and adaptability.
@@
Recent research has explored various control interfaces tailored to humanoid whole-body controller. These can be broadly categorized into dense and sparse control signals. Dense signals, such as teleoperation [1, 2, 3, 4, 5, 6], offline motion datasets [7, 8, 9, 10, 11, 12, 13], and video-based motion estimation [14, 15], provide rich trajectory-level information. In contrast, sparse signals such as high-level task commands and VR-based guidance [16, 17] offer minimal information and often lead to reduced motion quality. In this work, we focus on universal whole-body motion tracking, where the input is a reference motion sequence and the goal is to track it robustly and expressively using a single policy.
A widely adopted paradigm for learning motion tracking policies is the teacher-student framework. In this paradigm, a teacher policy is initially trained using full privileged observations to precisely track reference motions within simulation environments. Subsequently, a student policy is learned to imitate the teacher policy based solely on the partial observations available during deployment. Despite its effectiveness, existing implementations of this framework, particularly those utilizing simple MLP-based DAgger architectures [18] encounter three significant limitations. First, they frequently fail to preserve the diversity inherent in the original reference motions during the distillation process, leading to behavior that is averaged and less expressive. Second, constrained by their limited representational capacity, such models often exhibit poor generalization to unseen motion sequences. Third, the absence of global contextual information during training gives rise to issues such as orientation drift and broader inconsistencies in global behavior, which become particularly pronounced when the policies are deployed on real-world humanoid robots.
@@
Figure 2: An overview of UniTracker: In Stage 1, we train a teacher policy using oracle states via goal-conditioned reinforcement learning. In Stage 2, we distill the policy into a deployable form using a CVAE-based DAgger framework. In Stage 3, we introduce a fast adaptation module for handling challenging motion sequences, implemented using a residual decoder. The training dataset is derived from the AMASS dataset, filtered by PHC to remove physically infeasible motions.
To address the aforementioned limitations of existing teacher-student frameworks, we introduce UniTracker, a unified and expressive whole-body tracking framework that integrates a Conditional Variational Autoencoder (CVAE) [19] into the student policy architecture. By explicitly modeling a structured latent space conditioned on future motion references, UniTracker enables the policy to generate diverse and high-fidelity behaviors even under partial observations. From a probabilistic standpoint, the latent variable captures the inherent ambiguity in the mapping from observations to actions, allowing the policy to model a distribution over plausible motor behaviors rather than collapsing to a single deterministic output. This capability enhances motion expressiveness and significantly improves generalization across diverse and unseen movement patterns.
In addition, the CVAE-based framework also effectively addresses the challenge of missing global context—often manifested as orientation drift and other global inconsistencies during deployment. To this end, we employ task-aware feature modeling during training: the encoder is trained using privileged, globally informative observations to infer a structured latent representation, while a prior network is concurrently trained based only on the partial observations available at deployment time. The two distributions are aligned via a KL divergence objective. As a result, although the final deployed policy operates under partial observability, it benefits from a latent space informed by global context during training. This implicit incorporation of global information leads to more coherent and globally consistent behaviors in real-world settings.
While this CVAE-based universal policy demonstrates strong performance across a wide range of motions, it is neither necessary nor realistic to expect it to perfectly track all possible reference sequences—particularly those that are rare, highly dynamic, or lie far outside the training distribution. To accommodate such challenging cases, we introduce a fast adaptation phase that fine-tunes the universal policy in a task-specific manner. Leveraging the expressiveness and generality of the base policy, this adaptation process enables rapid specialization with minimal training time. Moreover, our framework supports both single-sequence adaptation and batch-mode adaptation, allowing for scalable refinement when dealing with multiple difficult motions. This final phase complements the universal policy by extending its practical applicability to edge cases, and highlights the modularity and flexibility of our overall three-stage training framework.
We extensively evaluate UniTracker in both simulated environments and real-world deployment scenarios. Experiments conducted on a 29-DoF Unitree G1 humanoid demonstrate that our policy is capable of tracking over 8,100 diverse motion sequences—including highly dynamic behaviors—using a single unified network. Compared to strong teacher-student baselines that do not incorporate CVAE-based modeling, UniTracker consistently achieves superior performance in terms of tracking accuracy, robustness, and generalization to unseen motions.
Our main contributions are summarized as follows: i)A three-stage training framework for universal whole-body tracking: We design a modular pipeline comprising a privileged teacher for data generation, a CVAE-based student policy for deployment under partial observations, and a lightweight fast adaptation phase for motion-specific fine-tuning. ii)Diversity-aware and gloabl context-integrated policy via CVAE modeling: We employ a Conditional Variational Autoencoder to capture motion diversity and encode global context, enabling expressive behaviors and reducing global inconsistencies by aligning a globally-informed encoder with a partial-observation prior. iii)Fast adaptation on top of a universal policy: We introduce a rapid fine-tuning mechanism that adapts the universal policy to challenging or out-of-distribution motions, supporting both single-sequence and batch-level adaptation with minimal overhead. iv)Extensive validation on real-world hardware: We demonstrate that UniTracker robustly tracks over 8,100 human motions on a 29-DoF humanoid using a single policy, outperforming strong teacher-student baselines in accuracy, robustness, and generalization.
2 Method
2.1 Problem Formulation
@@
We formulate the problem of humanoid robot whole-body motion tracking as a goal-conditioned reinforcement learning (RL) task, where a policy $\pi$ is trained to track reference motions at the whole-body level. The state $s_{t}$ comprises the robot’s proprioceptive information $s_{t}^{p}$ and the goal $s_{t}^{g}$ which specifies the target state for all body parts. The reward function $r_{t}=R(s_{t}^{p},s_{t}^{g})$, defined in terms of the agent’s proprioception and goal state, yields dense signals to guide policy optimization. To better focus on motion tracking at the whole-body level, we fix the wrist joints of our 29-degree-of-freedom (DoF) Unitree G1 robot[20], reducing the action space to 23 dimensions. The action $a_{t}\in\mathbb{R}^{23}$ specifies target joint positions, which are executed via a PD controller to actuate the robot. For policy optimization, we employ Proximal Policy Optimization (PPO) [21] to maximize the expected cumulative discounted reward $\mathbb{E}\left[\sum_{t=1}^{T}\gamma^{t-1}r_{t}\right]$.
The remainder of this section is organized as follows. Section 2.2 introduces the construction of a high-quality humanoid motion dataset for policy training. Section 2.3 describes the training of an oracle policy in simulation, which is designed both to maximize the expressiveness of motion tracking. Section 2.4 details the distillation of the oracle policy into a deployable student using a CVAE-based framework. Section 2.5 presents our fast adaptation strategy, which enables rapid fine-tuning based on the universal tracker. An overview of the proposed universal whole-body motion tracking framework is provided in Figure 2.
2.2 Humanoid Motion Dataset Curation
A large-scale humanoid motion dataset serves as fuel for training a universal motion tracker. Our dataset is primarily derived from the publicly available AMASS [22] dataset, filtered to exclude interactions and sequences shorter than 10 frames. This yields a final training set of 11,313 human motions, represented using the SMPL [23] parameters. The SMPL model parameterizes the human body through shape parameters $\beta\in\mathbb{R}^{10}$, pose parameters $\theta\in\mathbb{R}^{24\times 3}$ and root translation $p\in\mathbb{R}^{3}$. $S$ denotes the SMPL function, where $S(\beta,\theta,p):\beta,\theta,p\rightarrow\mathbb{R}^{6980\times 3}$ maps the SMPL parameters to the positions of vertices of a triangular human mesh.
To bridge the embodiment gap between SMPL human model and humanoid robots, we employ a two-stage retargeting approach inspired by H2O [4]. First, we carefully select 16 corresponding body links and optimize the shape parameter $\beta^{\prime}$ for humanoid robots by minimizing the distances between selected links in the rest pose. Second, leveraging the optimized $\beta^{\prime}$ alongside the original pose $\theta$ and translation $p$ from dataset, we perform gradient descent over the humanoid robot’s root translation, root orientation and joint positions to minimize the distances between selected links throughout the whole sequence. Additional regularization terms are added to avoid aggressive behaviors and ensure temporal smoothness.
2.3 Oracle Policy Training in Simulation
Reward Design. We formulate the reward $r_{t}$ as a weighted sum of three components: 1) task rewards for motion tracking, 2)regularization, and 3) penalty, detailed in Table 3. We apply curriculum learning to the regularization terms and penalty terms such that the policy could better focus on the motion tracking task itself and gradually take the penalty and regularization into account for more reasonable behaviors.
Early Termination and Reference State Initialization. In the early stage of training, the agent is prone to falling, resulting in the collection of invalid data that hinders effective learning. To address this issue, we follow prior works [24, 25] and introduce two early termination conditions: 1) orientation: the projected gravity on x or y axis exceeds 0.8; 2) tracking tolerance: the average link distance between the robot and reference motions is further than 0.5m. Proper task initialization is also crucial for RL training. We employ the Reference State Initialization [24, 9] framework, where the starting point of the reference motion is randomly sampled for the policy to track. The robot’s initial state, including the root position, orientation, linear and angular velocities, as well as joint positions and velocities, is then derived from the corresponding reference pose. This initialization strategy substantially enhances motion tracking training by enabling the policy to learn different motions phases in parallel, rather being constrained to a sequential learning process.
Domain Randomization. Domain randomization has been proved to be a critical source of robustness and generalization for successful sim-to-real transfer [11, 4, 6]. Existing domain randomization terms can be broadly categorized into two classes: asset property randomization and environmental dynamic randomization. Asset property randomization, focusing on parameters such as friction coefficients, center of mass offsets, and link masses, aims to prevent the policy from overfitting to a specific asset configuration. In contrast, environmental dynamic randomization, focusing on PD gains, applied torques, or other external perturbations (e.g., pushing the robot), introduces dynamic variations to simulate the environmental uncertainties encountered during real-world deployment. While previous works typically apply both forms of randomization concurrently throughout training, we propose decoupling them to better balance the tradeoff between motion tracking expressiveness and robustness. Hence, in the first stage, we exclusively randomize asset properties to maximize the policy’s ability to accurately track diverse motions. This approach not only builds an upper bound for subsequent policies but also yields a proxy agent capable of generating high-quality action signals in simulators.
2.4 Hierarchical Controller via Online Distillation
$\log$ $P(a_{t}|s_{t}^{p-d},s_{t}^{g-d})\geq E_{\varepsilon(z|s_{t}^{p-o},s_{t}^{g-o})}[\log D(a_{t}|s_{t}^{p-d},z_{t})$ (1)
$\rho(z_{t}|s_{t}^{p-d},s_{t}^{g-d})=\textit{N}$ $(\mu^{\rho}(s_{t}^{p-d},s_{t}^{g-d}),\sigma^{\rho}(s_{t}^{p-d},s_{t}^{g-d}))$ (2)
$\sigma^{\varepsilon}(s_{t}^{p-o},s_{t}^{g-o}))$
$L=L_{action}+\beta L_{KL}$ (3)
[…]
2.5 Fast Adaption on top of a Universal Policy
In this stage, we focus on motion references that the universal policy fails to reliably track. These motions typically exhibit two challenging characteristics: they lie far outside the training data distribution, and they involve highly dynamic behaviors that approach the physical limits of the humanoid robot.
To address these cases, we introduce a lightweight residual decoder $D(\widetilde{a_{t}}|{s_{t}}^{p-deploy},{s_{t}}^{g-deploy})$ appended to the universal motion tracking framework, to enable fast adaptation to challenging motions. During training, we initialize the model with the universal policy checkpoint and freeze all existing parameters. Only the residual decoder is fine-tuned using reinforcement learning, the reward is the same as Section 2.3. The input to the residual decoder is identical to that of the prior network. We directly use explicit motion references instead of latent representations for two main reasons. First, such out-of-distribution motions may not be adequately captured by the learned CVAE prior. Second, the goal at this stage is to quickly overfit to specific motion sequences rather than to maintain generalization. The final action applied to the environment is obtained by summing the output of the universal policy and the residual correction:
$a_{t}^{final}=a_{t}+\widetilde{a_{t}}$, (4)
where $a_{t}$ is the action predicted by the frozen universal policy, $\widetilde{a_{t}}$ is the residual generated by the residual decoder.
@@ 2.5 Fast Adaption on top of a Universal Policy
Train from Scratch 72.11 103.94 8.78 6.33 98.61 8.47 6.31
@@
Table 1: Baselines Comparison and Ablation Studies We compare our method with two baseline approaches and observe that it consistently outperforms both. Additionally, we conduct ablation studies on architectural design, KL residuals, KL coefficient, future window size, and latent dimensionality. Finally, we analyze the diversity of CVAE comparing the stochastic latent and the deterministic latent. The final configuration is selected based on the best overall performance observed across these settings.
3 Experimental Results
3.1 Experiment Setup